Estimate the minimum sample size needed to achieve the margin of error E equals= 0.015 for a​ 95% confidence interval. Round to the nearest integer.
Answer #1

We have that the ME = Zcritical * Sqrt(p * (1-p)/n).

Therefore n = (Zcritical/ME)2 * p * q

Zcritical at = 0.05 (2 tail) = 1.96 and ME is given as 0.015

When nothing is known, p = 1-p = 0.5

Therefore n = (1.96/0.015)2 * 0.5 * 0.5 = 4268.44 .

Rounding to the next Integer, n = 4269